You deposit $ 556 each month into an account earning 2.8 % interest compounded monthly for 6 years. How much total interest will you earn?

$_____________ (Round to the nearest DOLLAR)
Answer #1

sorry I read the question incorrectly. A = PMI Here PMT - $556 8o 0.028 n = 12 1286 t6 A = 596 (1 992) -,1 +6 28 2 4 35 36 TY A=P(1 + r) 556(1+0+0286 P =$556 ex/2.70209 = 0.028 0.028 C na 12 as total = 658 . you will earn $ (658-556)= $ 102 interest